OpenStudy (anonymous):

An expression is shown below: square root of 18 plus square root of 2 Which statement is true about the expression? It is rational and equal to 3. It is rational and equal to 4. It is irrational and equal to 3 multiplied by the square root of 2. It is irrational and equal to 4 multiplied by the square root of 2.

OpenStudy (anonymous):

@hhelpplzzzz @Love_Ranaa

OpenStudy (anonymous):

The functions f(x) and g(x) are shown below: f(x) = 0.05x g(x) = (0.05)x Which statement best describes the graph of f(x) and g(x)? The graph of g(x) will eventually exceed the graph of f(x). The graph of f(x) will eventually exceed the graph of g(x). The graphs will both have their y-intercept equal to 1. The graphs will both have their y-intercept equal to 0.05

OpenStudy (anonymous):

srry please help w. these guys i am on a time limit :)

OpenStudy (yolo_king):

\[\sqrt{18}= 3\sqrt{2}\]

OpenStudy (anonymous):

so c @YOLO_KING

OpenStudy (yolo_king):

\[\sqrt{2}=\sqrt{2}\]

OpenStudy (yolo_king):

Hold on

OpenStudy (anonymous):

ok @YOLO_KING

OpenStudy (yolo_king):

It is irrational and equal to 4 multiplied by the square root of 2. \[\sqrt{18} +\sqrt{2}=>3\sqrt{2}+\sqrt{2}\]

OpenStudy (yolo_king):

So that is for the first one did you understand?

OpenStudy (yolo_king):

IF so ask the next question by making a new post.

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

@YOLO_KING so it is the 1st one